OpenStudy (anonymous):

im doing addition of fractions and mixed numbers i don't get it! Can someone give me easy advice.?

OpenStudy (anonymous):

Do you have an example?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well if you can give me an example maybe I can help you

OpenStudy (anonymous):

Can you put your example on here?

OpenStudy (anonymous):

im also having trouble putting example im new on here .

OpenStudy (anonymous):

Ok so if it is a fraction for example writ 4/3.... Is that what you don't get?

OpenStudy (anonymous):

2/3 + 1/7 for example

OpenStudy (anonymous):

Ok so in order to add fractions you have to find a common denominator. what is something that both 3 and 7 go into?

OpenStudy (anonymous):

it can't

OpenStudy (anonymous):

Well 3(7)=21 so 21 is your common denominator. Now in order to get both your denominators to be 21 you have to adjust the fractions via multiplication: 2/3 3(7)=21 2(7)=14 2/3=14/21

OpenStudy (anonymous):

Now do the same thing to the other fraction so that the common denominator is 21

OpenStudy (anonymous):

The easy tip that mmbuckaroos is giving you is that you can mutliply the denominators to get a common denominator. This will cause you to have to simplify later in a lot of cases, but at least you are down to an answer at that point.

OpenStudy (anonymous):

i think im not getting the simplifying .

OpenStudy (anonymous):

Ok well we are not to the simplifying part yet... You want to multiply both the denominators to find a common denominator but whatever you do to the bottom of a fraction you have to do the same to the top of the fraction so for the fraction 2/3... Multiply the denominator (bottom) 3*7=21 Now you have to do the same thing to the top of the fraction 2*7=14 Your fraction now looks like this: 14/21

OpenStudy (anonymous):

Now do the same thing for the fraction 1/7....

OpenStudy (anonymous):

Multiply the denominator (bottom) 7*3=21 Now you have to do the same thing to the top of the fraction 1*3=3 Your fraction now looks like this: 3/21

OpenStudy (anonymous):

Your equation should look like this now: 14/21+3/21 Now you can simply add the numerators (top of the fraction) to get: 17/21

OpenStudy (anonymous):

17/21 can not be simplified so that is your final answer

OpenStudy (anonymous):

your steps are better than my instructors thanx. However i made a mistake it was 2/7+ 1/7 sorry

OpenStudy (anonymous):

Ok so that is even easier! You already have a common denominator All you have to do is add the numerators

OpenStudy (anonymous):

3/7 should be your answer

OpenStudy (anonymous):

so it stays 3/7 cuz it cant be simplified

OpenStudy (anonymous):

Yep:)

OpenStudy (anonymous):

how often do u help on here u ?

OpenStudy (anonymous):

Well Lately I have been on a lot trying to get some help myself... but problems like these are fun.

OpenStudy (anonymous):

i have always feared math but i have to finish my math to graduate from college. And im shy to get help w tutors at school.

OpenStudy (anonymous):

are u in los angeles?

OpenStudy (anonymous):

Understandable I'm not much of a math fan myself.. don't much like it at all really and no I'm not in Los Angeles... but what math are you taking if you are in college?

OpenStudy (anonymous):

Im actually a returning student and always put math off and now need to get it out of the way. I am taking math 20 which is basic.

OpenStudy (anonymous):

Nice! Hey better to get it done then never right?!? lol

OpenStudy (anonymous):

thanx so much !

OpenStudy (anonymous):

No problem anytime! I do online school so I basically have to teach myself cause my online teachers can only help you so much.....so I understand your frustrations. Just holler if you need anymore help:)

OpenStudy (anonymous):

well im going to work on the problems cuz homework is do tonight. When will u be on here again ?

OpenStudy (anonymous):

OH I'm on and off probably tomorrow... I have to catch up on some math HW so I'll be on here for as much help as I can get lol...

OpenStudy (anonymous):

are u a female or male just curious i like to know who im talking to.

OpenStudy (anonymous):

Female

OpenStudy (anonymous):

nice to meet u my name is Diana. I have a daughter she's math honors and won't really help me lol. u have been so kind. I will try and do my h.w. and if i have a question i will look for u tomorrow.

OpenStudy (anonymous):

Nice to meet you too:) Best of luck on your H.W. and sounds like a plan Stan;)